OpenStudy (anonymous):

Club a offers members fee of 70 plus monthly fee of 25. Club b offers members fee of 45 plus monthly fee of 30. Set up an equation for club membership cost. How many months will the cost of each club be the same. What will be total cost for each club from part b

hero (hero):

Hint: \(\large\frac{x}{2} + \frac{4x}{5} = \frac{2}{1}\)

hero (hero):

Multiply the third fraction by 2/2: \(\large\frac{x}{2} + \frac{4x}{5} = \frac{4}{2}\)

hero (hero):

Subract x/2 from both sides to get: \(\large\frac{4x}{5} = \frac{4-x}{2}\)

hero (hero):

Cross Multiply to get 2(4x) = 5(4-x)

hero (hero):

Now continue solving for x

hero (hero):

You should get \(x = \frac{20}{13}\)

OpenStudy (anonymous):

Thanks

hero (hero):

@izzyboe2, please change the question back to what it was and re-post the original one. You have to "close this question" if you want to post a new question. Otherwise, you make my steps look bad.

hero (hero):

@izzyboe2 , you have to change the question back to what it was before, otherwise students who are searching for the previous question won't be able to find it.

hero (hero):

In order to post another question, you have to "close" this one

OpenStudy (anonymous):

Ok